When an elastic material with Youngs modulus Y is subjected to stretching stress S, elastic energy stored per unit volume of the material is  


        (a)    YS/2


        (b)    S2Y/2


        (c)    S2/2Y        


        (d)    S/2Y


Grade:upto college level

2 Answers

Yuvraj
10 years ago
(c) As the elastic energy stored=1/2(Stress)(strain) and strain=Stress/Y
Soumyajeet Das
5 years ago
U= ½ X Stress X Strain X Volume
U/Volume = ½ X Stress X Strain
U/V= ½ X F/A X dL/L  
[Since, F/A =Y X dL/L]
Therefore,
U/V= ½ X Y X (dL/L)^2
U/V=1/2 X Y X (S)^2
       =Y(S)^2/2
Option B